A silver wire that is 2.5 mm in diameter and 37 cm long carries a current of 169 mA. How many electrons per second pass a given cross-section of the wire?


1
Expert's answer
2021-05-21T10:43:12-0400

By the definition of the current we have:


"I=\dfrac{Q}{t},""Q=It."

From the other hand,


"Q=ne."

Therefore, we can write:


"It=ne,""\dfrac{n}{t}=\dfrac{I}{e},""\dfrac{n}{t}=\dfrac{169\cdot10^{-3}\ A}{1.6\cdot10^{-19}\ C}=1.06\cdot10^{18}\ \dfrac{electrons}{s}."